OpenStudy (anonymous):

A carpenter is building a rectangular room with a fixed perimeter of 376 ft. What dimensions would yield the maximum area? What is the maximum area?

OpenStudy (anonymous):

2l+2w=376 lw=a these are the 2 equations u need differentiate the area in terms of length so u need to solve for w in the first equation then substitute into second equation. after that differentiate the equation u substituted to and set it equal to 0

OpenStudy (anonymous):

l = w = 94

OpenStudy (anonymous):

376/4=94 94^2=8836

OpenStudy (anonymous):

also when perimeter is limited and u have to choose between a rectangle and a square, a square holds more area.